# Hidden in Plain Sight II: Characterizing the luminous companion to $\kappa$ Velorum with VLTI/GRAVITY

### v1

@@ 7.1 Other VLTI Targets
Support for this work was provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51588.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. S.K. acknowledges support from an European Research Council (ERC) Consolidator Grant (Grant Agreement ID 101003096), as well as an STFC Small Award (ST/Y002695/1). The VLTI GRAVITY observations were collected at the European Organisation for Astronomical Research in the Southern Hemisphere under ESO programme 114.274C. The UVES analysis is based on observations collected at the European Organisation for Astronomical Research in the Southern Hemisphere under ESO programme 194.C-0833. This research has made use of the SIMBAD database, operated at CDS, Strasbourg, France. This research has made use of the VizieR catalogue access tool, CDS, Strasbourg, France.

### v2

@@ VII.1 Other VLTI Targets
We thank the anonymous reviewer for their useful comments that improved the quality of this manuscript. DMR and TAT thank Chris Kochanek and Kris Stanek for helpful conversations. TAT thanks Keith Smith for motivating questions about the Kappa Velorum system. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). Support for this work was provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51588.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. S.K. acknowledges support from an European Research Council (ERC) Consolidator Grant (Grant Agreement ID 101003096), as well as an STFC Small Award (ST/Y002695/1). The VLTI/GRAVITY observations were collected at the European Organisation for Astronomical Research in the Southern Hemisphere under ESO programme 114.274C. The UVES analysis is based on observations collected at the European Organisation for Astronomical Research in the Southern Hemisphere under ESO programme 194.C-0833. This research has made use of the SIMBAD database, operated at CDS, Strasbourg, France. This research has made use of the VizieR catalogue access tool, CDS, Strasbourg, France.